OpenStudy (anonymous):

Factor x^3 - 2x^2 - 16x + 32 = 0, given that 2 is a zero. ANSWERS BELOW**

OpenStudy (anonymous):

A.(x -2)(x^2 -16) = 0 B.(x + 2)(x -4)(x + 4) = 0 C.(x -2)(x -4)(x + 4) = 0

OpenStudy (anonymous):

HELP PLEASE @Cydney_Morgan

OpenStudy (cydney_morgan):

Give me a moment :)

OpenStudy (anonymous):

thanks so much!:)

OpenStudy (fibonaccichick666):

So if 2 is a zero, what does that mean?

OpenStudy (anonymous):

idfk???!

OpenStudy (fibonaccichick666):

...language.

OpenStudy (anonymous):

freaking?

OpenStudy (fibonaccichick666):

anyways, if you have (x-5)(x-7)=0 What are the zeroes?

OpenStudy (anonymous):

dude i really dont know?

OpenStudy (fibonaccichick666):

let's get a little more basic then, what is a zero?

OpenStudy (anonymous):

a number duh

OpenStudy (fibonaccichick666):

... a little more than a number.